OpenStudy (anonymous):

Using homogenous equation , solve , ( x^2 + y^2 ) dx + xydy = 0

OpenStudy (shubhamsrg):

divide both sides by x^2 and see if it helps your case.

OpenStudy (anonymous):

Skylar needs $56 to buy a new video game. If he doubles the amount of money he has now, he will have enough to buy the game with $8 left over. How much money does Skylar have right now

OpenStudy (anonymous):

sagutin ko hahaha

OpenStudy (anonymous):

aha bat nandito ka

OpenStudy (anonymous):

nandun na ko sa may ln|x| + 1/2| 1 + 2v^2 |

OpenStudy (anonymous):

Guys, i arrived at ln|x| + 1/2 | 1+2v^2 | = 0 , i can't simplify it

OpenStudy (anonymous):

let y=vx dy=vdx + xdv substitute: \[(x^{2}+v^2x^2)dx+vx^2(vdx+xdv)=0\] \[x^2dx+v^2x^2dx+v^2x^2dx+v^3dv=0\] \[x^2dx+2v^2x^2dx+vx^3dv=0\] \[x^2(1+2v^2)dx+vx^3dv=0 ] \frac{ 1 }{ x^3(1+2v^2) }\] \[\int\limits_{}^{}\frac{ dx }{ x }+\int\limits_{}^{}\frac{ vdv }{ 1+2v^2 }=0\] \[4(lnx+\frac{ 1 }{ 4 }\ln \left| 1+2v^2 \right|=c\] \[4lnx+\ln \left| 1+2v^2 \right|=4c\] \[lnx^{4}+\ln \left| 1+2v^2 \right|=4c\] \[\ln \left\{ x^4\left( 1+2v^2 \right) \right\}=4c\] let v=\[\frac{ y }{ x }\]

OpenStudy (anonymous):

di pa tapos

OpenStudy (anonymous):

substitute mo value ng v sa \[\ln{x^4(1+2v^2)}=4c\]

OpenStudy (anonymous):

bukas nlng haha